OpenStudy (anonymous):

I need help on this question!!! f(x)=(e^x)/(e^x+2)(x+3) find the derivative f'(x)=_____ Please help me!!

zepdrix (zepdrix):

It's hard to read it properly with the way you formatted it. Is this correct?\[\Large f(x)\quad=\quad \frac{e^x}{(e^x+2)(x+3)}\]

OpenStudy (anonymous):

that is correct!

zepdrix (zepdrix):

\[\Large f'(x)=\frac{\color{royalblue}{\left[e^x\right]'}(e^x+2)(x+3)-e^x\color{royalblue}{\left[(e^x+2)(x+3)\right]'}}{\left[(e^x+2)(x+3)\right]^2}\]

zepdrix (zepdrix):

Oh boy this is going to be a doozy. So there is our quotient rule setup. We'll need to take the derivatives of the blue portions.

zepdrix (zepdrix):

Looks like we'll need to apply the product rule to the second blue part. Any confusion on the quotient rule setup? :o

OpenStudy (anonymous):

no I am pretty comfortable with the quotient rule

OpenStudy (loser66):

jealous !! he makes the explanation clear with color.

OpenStudy (anonymous):

i have the correct set up, for some reason i am not carrying my math out right

OpenStudy (anonymous):

nice use of \(\color{blue}{color}\)

OpenStudy (anonymous):

i still do not understand even with the color can someone explain it to me please haha

zepdrix (zepdrix):

Ah sorry I ran off for a sec there :d

zepdrix (zepdrix):

So what does the derivative of the first blue part give you? e^x comes out of that one, right? :o

zepdrix (zepdrix):

\[\Large f'(x)=\frac{\color{orangered}{\left[e^x\right]}(e^x+2)(x+3)-e^x\color{royalblue}{\left[(e^x+2)(x+3)\right]'}}{\left[(e^x+2)(x+3)\right]^2}\]

OpenStudy (anonymous):

i believe that is correct

zepdrix (zepdrix):

So I guess the tricky part is the other blue part, product rule time!\[\Large \left[(e^x+2)(x+3)\right]'=\color{teal}{\left(e^x+2\right)'}(x+3)+(e^x+2)\color{teal}{\left(x+3\right)'}\]

OpenStudy (anonymous):

hmm would that be e^x(x+3) + (e^x +2)(1)

zepdrix (zepdrix):

Mmmmm yah looks good!

zepdrix (zepdrix):

\[\Large f'(x)=\frac{\color{orangered}{\left[e^x\right]}(e^x+2)(x+3)-e^x\color{orangered}{\left[e^x(x+3)+(e^x+2)\right]}}{\left[(e^x+2)(x+3)\right]^2}\]

zepdrix (zepdrix):

So from thereeeeee, it's just some... messy... simplification :\

zepdrix (zepdrix):

imma go make a sammich, simplifyyyy \c:/

OpenStudy (anonymous):

haha alright so the squared on the bottom gets cancelled because of the ones in the numerator

zepdrix (zepdrix):

The square? yah that sounds right, leaving us with a 1 exponent down there.

zepdrix (zepdrix):

Err wait wait :p

zepdrix (zepdrix):

We don't have an \(\Large (e^x+2)(x+3)\) in the second term :( Hmm

OpenStudy (anonymous):

right then it would be -e^x/ (e^x+2)(x+3)

OpenStudy (anonymous):

would it?

OpenStudy (anonymous):

how could i simplify this further??

zepdrix (zepdrix):

I dunnoooo :( too much multiplication.. brain.. hurts.

OpenStudy (anonymous):

is the answer this -e^x/ (e^x+2)(x+3) because i plugged that in and i was told it was wrong

OpenStudy (anonymous):

haha believe me my brain hurts too i have been on this problem for the last two hours

zepdrix (zepdrix):

bahh i dunno +_+ ill multiply it out after i finish this slice of pizza

OpenStudy (anonymous):

thank you so much i really appreciate it!

zepdrix (zepdrix):

Mmm it looks like this is not going to simplify nicely, you should try to put it all into your answer. :\

zepdrix (zepdrix):

I'm not sure how you got (e^x+2)(x+3) in your denominator... Hmm

OpenStudy (anonymous):

good point im not sure either haha im so confused with this problem i understand the rules but for some reason this problem is getting to me

OpenStudy (anonymous):

you are an absolute genius thank you so very much!! you saved me big time!!!!!!!!